OpenStudy (elaneemarie):

which pairs of ratios form a proportion? a-10/20 and 30/40 b-3/7 and 39/78 c-32/80 and 4/10 d-45/56 and 5/7

OpenStudy (radar):

Divide the pairs (ratios) and the pair that results in an equality is the proportion. like. a 10/20 = 1/2 and 30/40 = 3/4 Not the same. a. is out. just go down the list.

OpenStudy (radar):

to check out b divide 3/7 getting .428 and 39/78=1/2 or .5 close, but not equal rejecting b. ....etc.....

OpenStudy (triciaal):

why is this still open?